# An open source Multi-Agent Deep Reinforcement Learning Routing Simulator for satellite networks

### v1

@@ Abstract
This paper introduces an open source simulator for packet routing in Low Earth Orbit Satellite Constellations (LSatCs) considering the dynamic system uncertainties. The simulator, implemented in Python, supports traditional Dijkstra’s based routing as well as more advanced learning solutions, specifically Q-Routing and Multi-Agent Deep Reinforcement Learning (MA-DRL) from our previous work. It uses an event-based approach with the SimPy module to accurately simulate packet creation, routing and queuing, providing real-time tracking of queues and latency. The simulator is highly configurable, allowing adjustments in routing policies, traffic, ground and space layer topologies, communication parameters, and learning hyperparameters. Key features include the ability to visualize system motion and track packet paths. Results highlight significant improvements in end-to-end (E2E) latency using Reinforcement Learning (RL)-based routing policies compared to traditional methods. The source code, the documentation and a Jupyter notebook with post-processing results and analysis are available on GitHub.
@@ 1 Introduction
Efficient routing in LSatCs is critical for global connectivity in 6G networks. This requires addressing multiple challenges, including the partial knowledge of the network at the satellites and their continuous movement, and the time-varying sources of uncertainty in the system, such as traffic, communication links, or communication buffers [lozanocuadra2024continual]. Traditional routing algorithms are inadequate to address these problems: They either lack adaptability to network changes or congest the network with feedback messages. To overcome these challenges, new algorithms must be developed, some of them RL-based, which need to be accompanied by a robust framework. Python is the best environment for developing RL-based algorithms due to its extensive libraries for machine learning, such as Keras-TensorFlow, PyTorch, NumPy, and Pandas. This paper introduces an open source MA-DRL Routing Simulator for satellite networks built in Python, where these designed algorithms can be implemented and tested. The simulator currently supports various routing algorithms, including some Dijkstra’s [dijkstra1959] shortest path-based and those from our recent works: (1) The Q-Routing with Q-tables for distributed routing decisions [soret2023qlearning], and (2) the MA-DRL first proposed in [lozanocuadra2024multiagent], which was then further tested and extended to continual learning with Satellite Federated Learning (SFL) in [lozanocuadra2024continual]. The source code, a Jupyter notebook with some post-processing results and analysis, and the documentation for the MA-DRL Routing Simulator are available on GitHub [lozano2024madrl].
Figure 1: Kepler constellation deployed and their corresponding ISLs established following the Greedy matching with 18 active gateways over the population maps [CIESIN], where the green tone depends on the population density. Each satellite´s colour is a different orbital plane.
@@ 2 Simulator architecture
The environment is an event-based Python simulator using the SimPy module, chosen for its effectiveness in discrete event modeling [zinoviev2024discrete]. Time in the simulator progresses by jumping from one scheduled event to the next, rather than continuously. Each action, including creating, routing, and queueing of individual data packets, is explicitly simulated as a SimPy event, providing accurate real-time tracking of queues and latency. Packets are unique entities (objects) existing from creation at a generating gateway until they arrive at the destination gateway. The transmission time is calculated based on the packet size and current link rates, propagation time based on the exact distance between transmitter and receiver at transmission time, and queue time based on the time a packet spends in the queue. This detailed level of simulation is essential for representing the states and computing the rewards of the environment in our RL-based routing algorithms.
The simulator emulates a realistic scenario where ground gateways gather the nearby terrestrial traffic that is assumed to be generated by mobile users and distribute that to each other gateway equally through a LSatC, integrating space and ground segments into the communication network, as shown in Fig. 1. Computationally, this is built as a time-variant dynamic graph $\mathcal{G}_{t}(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}$, representing satellites and gateways, and edges $\mathcal{E}$, representing the transmission links between them, which can be either ISL or ground-to-satellite link (GSL), implemented as Radio Frequency (RF) or Free Space Optical (FSO).
Space segment. The satellite constellation consists of $N$ satellites evenly distributed across $O$ orbital planes. Each satellite functions as a router and learning agent for RL-based solutions. Satellites are positioned at specific and configurable altitudes, longitudes, and orbit inclinations, moving according to orbital mechanics and Earth’s rotation [lozanocuadra2024continual]. Satellites move periodically rather than continuously. After a fixed time interval, each satellite jumps to its new position as if it had moved continuously during that period. This periodic movement impacts latency calculations by updating transmission and propagation times at each position update. Each satellite has one antenna for GSL and four for ISL (two for inter-plane and two for intra-plane links). Selecting the best ISL is a dynamic matching problem and consists of establishing the best ISLs among satellites. Links are bidirectional, and the network is reconfigured as satellites move, i.e, $\mathcal{G}_{t}$ is built again maintaining previous queue states.
@@
Figure 3: MA-DRL ´s exploitation phase congestion test for all routes output with 8 active gateways.
@@ 4 Setup and general settings
To run the simulator, the user needs Python 3.9. The simulator is multi-platform and has been tested on Windows, Linux, and Mac systems. The user can install the required packages listed in the requirements.txt file using pip. It is advisable to create a virtual environment or an Anaconda environment for better management.
The simulator is highly configurable, allowing users to adjust various parameters to suit their specific needs, as illustrated in Fig. 2. Key configurable parameters include: (a) Routing Policy, including the shortest path-based, where the Data Rate, Slant Range or Hop can be set as weights, and the RL-based options; (b) Ground Layer settings, such as the number and locations of active gateways, as well as traffic generation $\ell$; (c) Space Layer parameters, which cover constellation design (configurable elements include the number of orbital planes and its inclination angle, satellites per plane, and the choice between Walker delta and Walker star designs), ISL matching (Greedy or Markovian [Leyva-Mayorga2021]), and orbital motion (which can be sped up or slowed down); (d) Learning Hyperparameters, including rewards and penalties, exploration $\epsilon$, learning $\alpha$ and gamma $\gamma$ rates, state preprocessing, and training modes (Import pre-trained models for RL-based policies and choosing between online and offline phases for MA-DRL); and (e) Communication Setup, such as physical constants, uplink and downlink parameters, and packet size. Additionally, the user can configure the simulator to plot the environment every time the constellation moves to visualize system motion, as in Fig. 1, and to plot the path of each delivered packet over this to track its journey through the network.
@@ 5 Results and analysis
There are up to $18$ defined active (transmitting and receiving) gateways distributed across the Earth, mainly following KSAT’s deployment [^1] https://www.ksat.no/services/ground-station-services/, however, more gateways can be configured. Moreover, four real constellations are implemented: (1) Kepler constellation design, with $O=7$ orbital planes at heights $h=600$ km and $N=20$ satellites per orbital plane, as illustrated in Fig. 1; (2) Iridium Next constellation, with $O=6$, $h=780$ km and $N=11$; (3) the OneWeb constellation, with $O=36$, $h=1200$ km and $N=18$; and (4) Starlink orbital shell at $h=550$ km, with $O=72$ and $N=22$. The three first constellations follow a Walker star architecture, while the Starlink shell follows a Walker delta architecture [leyva2022ngso]. Moreover, two additional artificial constellations are implemented for testing. Additionally, two ISL matching algorithms are implemented: (1) The one Markovian solution in [Leyva-Mayorga2021] and (2) a Greedy approach, where each satellite connects with immediate neighbors within its plane and closest counterparts in adjacent planes in both East and West directions, optimizing for latency and data rates, as shown in Fig. 1.
@@
Every time a simulation ends, it outputs automatically several results and analysis of what happened in figures and files (Fig. 2). Within the figures, a map like Fig. 1 with the system model information is saved. An update of this figure is also saved as the constellation moves if desired. Then, a congestion test per route and for all routes between gateways is done, in order to see what nodes and edges did the packets went through, as shown in Fig. 3. If one of the RL-based routing policies is chosen, one figure with the exploration rate $\epsilon$ and training stamps and another one with the received rewards (Fig. 4) are also saved. Other figures saved are related to the average E2E latency vs time vs $\epsilon$, similar to Fig. 5, but with just one routing policy, and to the queue lengths. On the other hand, the output files include several .csv with extensive information about each packet´s path and its latency, rewards, exploration rates, training stamps, hyper-parameters and a .txt log-file, that saves everything that happened during the simulation and gives some statistics like like the latency broken down by average queue, transmission and propagation times, packets delivered vs stuck and/or lost, most used links, etc. Lastly, if either the MA-DRL or the Q-Routing algorithm was chosen for routing, the trained DNNs (57Kb for the Q-Network and 27 Kb for the Q-Target) or Q-Tables (21Kb for 8 active gateways) are saved, respectively.
@@
Figure 6: Average E2E latency over an orbital period. The fluctuations are given by the movement of the satellites and the consequent changes in the followed routes by the packets.
@@
In the Jupyter notebook, we conduct further post-processing analysis and explore more complex results. A comparison between the Shortest path routing policy, Q-Routing [soret2023qlearning] and MA-DRL [lozanocuadra2024multiagent, lozanocuadra2024continual] at their offline phase is shown in Fig. 5. Additionally, a dynamic comparison of these policies at their online phase is shown in Fig. 6, where the constellation has moved to complete one orbital period in $96$ minutes, where satellite positions were updated at intervals of $15$ seconds. Notably, even with only partial knowledge of the constellation, MA-DRL consistently maintains the baseline latency obtained with the Shortest path policy, which has full knowledge of the constellation. Moreover, we elaborate on the comparison of the four architectures in Fig. 7, where the distribution of the E2E latency is depicted in a box plot when the Shortest path is applied among one orbital period too. We observe that Kepler and Starlink obtain the smallest average latency, although the latter presents more outliers. This figure helps to illustrate the behavior of the constellations and highlights the usage of the simulator to test different constellation architectures.
@@ 6 Conclusion and Future work
Future directions include: (1) Developing a SFL framework to enable aggregation during the online phase of MA-DRL rather than implementing it as a post-processing analysis; (2) implementing a two-tier mesh network for UE-satellite-UE communications, enabling ground moving users to connect directly to satellites without the need for gateways; (3) increase the complexity of satellites with regenerative capabilities; and (4) implementing different types of traffic with splittable flows.

### v2

@@ Abstract
This paper introduces an open source simulator for packet routing in Low Earth Orbit Satellite Constellations (LSatCs). The simulator, implemented in Python, supports traditional Dijkstra’s based routing as well as more advanced learning solutions based on Q-Routing and Multi-Agent Deep Reinforcement Learning (MA-DRL) from our previous work. It uses an event-based approach with the SimPy module to accurately simulate packet creation, routing and queuing, providing real-time tracking of queues and latency. The simulator is highly configurable, allowing adjustments in routing policies, traffic, ground and space segment topologies, communication parameters, and learning hyperparameters. Key features include the ability to visualize system motion and track packet paths while considering the inherent uncertainties of such a dynamic system. Results highlight significant improvements in end-to-end (E2E) latency using Reinforcement Learning (RL)-based routing policies compared to traditional methods. The source code, the documentation and a Jupyter notebook with post-processing results and analysis are available on GitHub.
@@ 1 Introduction
Efficient routing in LSatCs is critical for global connectivity in 6G networks. This requires addressing multiple challenges, including the partial knowledge of the network at the satellites and their continuous movement, and the time-varying sources of uncertainty in the system, such as traffic, communication links, or communication buffers [lozanocuadra2024continual]. Traditional routing algorithms are inadequate to address these problems: They either lack adaptability to network changes or congest the network with feedback messages. To overcome these challenges, new algorithms must be developed, some of them RL-based, which need to be accompanied by a robust framework. Python is the best environment for developing RL-based algorithms due to its extensive libraries for machine learning, such as Keras-TensorFlow, PyTorch, NumPy, and Pandas. This paper introduces an open source MA-DRL Routing Simulator for satellite networks built in Python, where these designed algorithms can be implemented and tested. The simulator supports various routing algorithms, including some Dijkstra’s [dijkstra1959] shortest path-based and those from our recent works: (1) The Q-Routing with Q-tables for distributed routing decisions [soret2023qlearning], and (2) the MA-DRL first proposed in [lozanocuadra2024multiagent], which was then further tested and extended to continual learning with Satellite Federated Learning (SFL) in [lozanocuadra2024continual]. The source code, a Jupyter notebook with some post-processing results and analysis, and the documentation for the MA-DRL Routing Simulator are available on GitHub [lozano2024madrl].
Figure 1: Kepler constellation deployed and their corresponding ISLs established following the Greedy matching with 18 active gateways over the population maps [CIESIN], where the green tone depends on the population density. Each satellite’s colour is a different orbital plane.
@@ 2 Simulator architecture
The event-based simulation environment was developed in Python using the SimPy module, chosen for its effectiveness in discrete event modeling [zinoviev2024discrete]. Time in the simulator progresses by jumping from one scheduled event to the next, rather than continuously. Each action, including creating, routing, and queueing of individual data packets, is explicitly simulated as a SimPy event, providing accurate real-time tracking of queues and latency. Packets are unique entities (objects) existing from creation at a generating gateway until they arrive at the destination gateway. The transmission time is calculated based on the packet size and current link rates, propagation time based on the exact distance between transmitter and receiver at transmission time, and queue time based on the time a packet spends in the queue. This detailed level of simulation is essential for representing the states and computing the rewards of the environment in our RL-based routing algorithms.
The simulator emulates a realistic scenario where ground gateways gather the nearby terrestrial traffic that is assumed to be generated by mobile users and distribute that to each other gateway equally through a LSatC, integrating space and ground segments into the communication network, as shown in Fig. 1. The environment is built as a time-variant dynamic graph $\mathcal{G}_{t}(\mathcal{N},\mathcal{E})$ with nodes $\mathcal{N}$, representing satellites and gateways, and edges $\mathcal{E}$, representing the transmission links between them, which can be either ISL or ground-to-satellite link (GSL), implemented as Radio Frequency (RF) or Free Space Optical (FSO).
Space segment. The satellite constellation consists of $N$ satellites evenly distributed across $O$ orbital planes. Each satellite functions as a router and learning agent for RL-based solutions. Satellites are positioned at specific and configurable altitudes, longitudes, and orbit inclinations, moving according to orbital mechanics and Earth’s rotation [lozanocuadra2024continual]. Satellites move periodically, at the beginning of each time interval, rather than continuously. After a fixed time interval, each satellite is placed in the exact position that it would reach if it had moved continuously during that period. This periodic movement impacts latency calculations by updating transmission and propagation times at each position update. Each satellite has one antenna for GSL and four for ISL (two for inter-plane and two for intra-plane links). Selecting the best ISL is a dynamic matching problem and consists of establishing the best ISLs among satellites. Links are bidirectional, and the network is reconfigured as satellites move, i.e, $\mathcal{G}_{t}$ is built again maintaining previous queue states.
@@
Figure 3: MA-DRL’s exploitation phase congestion test for all routes output with 8 active gateways.
@@ 4 Setup and general settings
The simulator, running in Python 3.9, is multi-platform and has been tested on Windows, Linux, and Mac systems. The user can install the required packages listed in the requirements.txt file using pip. It is advisable to create a virtual environment or an Anaconda environment for better management.
The simulator is highly configurable, allowing users to adjust various parameters to suit their specific needs, as illustrated in Fig. 2. Key configurable parameters include: (a) Routing Policy, including the shortest path-based, where the Data Rate, Slant Range or Hop can be set as weights, and the RL-based options; (b) Ground Segment settings, such as the number and locations of active gateways, as well as traffic generation $\ell$; (c) Space Segment parameters, which cover constellation design (configurable elements include the number of orbital planes and its inclination angle, satellites per plane, and the choice between Walker delta and Walker star designs), ISL matching (Greedy or Markovian [Leyva-Mayorga2021]), and orbital motion (which can be sped up or slowed down); (d) Learning Hyperparameters, including rewards and penalties, exploration $\epsilon$, learning $\alpha$ and gamma $\gamma$ rates, state preprocessing, and training modes (Import pre-trained models for RL-based policies and choosing between online and offline phases for MA-DRL); and (e) Communication Setup, such as physical constants, uplink and downlink parameters, and packet size. Additionally, the user can configure the simulator to plot the environment every time the constellation moves to visualize system motion, as in Fig. 1, and to plot the path of each delivered packet over this to track its journey through the network.
@@ 5 Results and analysis
The default ground segment has up to $18$ active (transmitting and receiving) gateways distributed across the Earth, mainly following KSAT’s deployment [^1] https://www.ksat.no/services/ground-station-services/, but more gateways can be added easily.
@@
Moreover, four real constellations are implemented: (1) Kepler constellation design, with $O=7$ orbital planes at heights $h=600$ km and $N=20$ satellites per orbital plane, as illustrated in Fig. 1; (2) Iridium Next constellation, with $O=6$, $h=780$ km and $N=11$; (3) the OneWeb constellation, with $O=36$, $h=1200$ km and $N=18$; and (4) Starlink orbital shell at $h=550$ km, with $O=72$ and $N=22$. The three first constellations follow a Walker star architecture, while the Starlink shell follows a Walker delta architecture [leyva2022ngso]. Moreover, two additional artificial constellations are implemented for testing. Additionally, two ISL matching algorithms are implemented: (1) The Markovian solution proposed in [Leyva-Mayorga2021] and (2) a Greedy approach, where each satellite connects with immediate neighbors within its plane and closest counterparts in adjacent planes in both East and West directions, optimizing for latency and data rates, as shown in Fig. 1.
@@
When a simulation ends, it automatically outputs a set of results in the form of figures and text files (Fig. 2). Within the figures, a map like Fig. 1 with the system model information is saved. An update of this figure is also saved as the constellation moves if desired. Then, a congestion test per route and for all routes between gateways is done, in order to see what nodes and edges did the packets went through, as shown in Fig. 3. If one of the RL-based routing policies is chosen, one figure with the exploration rate $\epsilon$ and training stamps and another one with the received rewards (Fig. 4) are also saved. Other figures saved are related to the average E2E latency vs time vs $\epsilon$, similar to Fig. 5, but with just one routing policy, and to the queue lengths. On the other hand, the output files include several .csv with extensive information about each packet’s path and its latency, rewards, exploration rates, training stamps, hyper-parameters and a .txt log-file, that saves everything that happened during the simulation and gives some statistics like like the latency broken down by average queue, transmission and propagation times, packets delivered vs stuck and/or lost, most used links, etc. Lastly, if either the MA-DRL or the Q-Routing algorithm was chosen for routing, the trained DNNs (57Kb for the Q-Network and 27 Kb for the Q-Target) or Q-Tables (21Kb for 8 active gateways) are saved, respectively.
@@
Figure 6: Average E2E latency over an orbital period. The fluctuations are given by the movement of the satellites and the resulting changes in the routed followed by the packets.
@@
In the Jupyter notebook, we conduct further post-processing analysis and explore more complex results. A comparison between the Shortest path routing policy, Q-Routing [soret2023qlearning] and MA-DRL [lozanocuadra2024multiagent, lozanocuadra2024continual] at their offline phase is shown in Fig. 5. Additionally, a dynamic comparison of these policies at their online phase is shown in Fig. 6, where the constellation has moved to complete one orbital period in $96$ minutes, with the satellite positions being updated at intervals of $15$ seconds. Notably, even with only partial knowledge of the constellation, MA-DRL consistently maintains the baseline latency obtained with the Shortest path policy, which has full knowledge of the constellation. Moreover, we elaborate on the comparison of the four architectures in Fig. 7, where the distribution of the E2E latency is depicted in a box plot when the Shortest path is applied among one orbital period too. We observe that Kepler and Starlink obtain the smallest average latency, although the latter presents more outliers. This figure helps to illustrate the behavior of the constellations and highlights the usage of the simulator to test different constellation architectures.
@@ 6 Conclusion and Future work
Future directions include: (1) Developing an SFL framework to enable aggregation during the online phase of MA-DRL rather than implementing it as a post-processing analysis; (2) implementing a two-tier mesh network for UE-satellite-UE communications, enabling ground moving users to connect directly to satellites without the need for gateways; (3) increase the complexity of satellites with regenerative capabilities; and (4) implementing different types of traffic with splittable flows.